Assume that a stock price increased from 20 to 24 in 3 months. What is the implied annual volatility in the context of one step binomial trees?

a-0,205

b-0,425

c-0,365

d-0,305
Answer #1

S0*exp(volatility*sqrt(t))=Su
=>20*exp(volatility*sqrt(3/12))=24
=>volatility*sqrt(3/12)=ln(24/20)
=>volatility=ln(24/20)*1/sqrt(3/12)
=>volatility=36.4643%=0.364643